• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yoomin

    Log entry time 07:49:02 on February 11, 2009

    Entry number 260461

    keyword=Owl Shift Summary

    Owl Shift Summary Feb 11, 2009

    00:00 BCM calibration run is in progress. Q3 oscillating : 0.867~0.87015(set value)~0.8735 GeV 01:15 Give up BCM calibration due to lack of time 01:35 Check beam with optics/BeO and ref. cell. 01:50 Start BB optics with Xin, 15uA, 4x6 rastered by MCC unit. 03:00 Ask for controlled access, Brad is coming in to put BB sieve OUT. 03:10 Controlled access 03:25 Controlled access end, BB SIEVE IS OUT. 03:40 Q2/Q3 cycled to ramp up, but just got back to original setting. 04:00 Take two runs of optics target with 5uA beam, 1286 with raster on (4x7 MCC) and 1288 with raster off 04:40 Back to H2 ref cell and taking BB optics runs 05:00 Take data for tracking efficiency study 07:00 Ask for controlled access as planned. Target move to empty. 07:10 Set L-HRS magnet fields to 0 to prepare polarity change. 07:45 Under controlled access.

    Run List -BB optics with H2 reference cell Run Time #Events Comment 1281 02:01-02:24 1M 15uA, raster 4x6 MCC 1282 02:25-02:47 1M 15uA 1283 02:47-03:05 1M 15uA 1289 04:45-05:12 1.5M 15uA 1291 05:27-05:54 1M 12uA 1292 05:56-06:21 1M 8uA 1293 06:22-06:42 550k 4uA, BB HV trip 1294 06:47-07:03 780k 2uA, controlled access -BB optics with multi carbon foil w/ BeO target Run Time #Events Comment 1286 04:21-04:26 1M raster on 4x7 MCC, 5uA 1288 04:05-04:14 1M raster off, 5uA